# Rakuten Data Release: A Large-Scale and Long-Term Reviews Corpus for Hotel Domain

### v2

@@ Abstract
This paper presents a large-scale corpus of Rakuten Travel Reviews. Our collection contains 7.29 million customer reviews for 16 years, ranging from 2009 to 2024. Each record in the dataset contains the review text, its response from an accommodation, an anonymized reviewer ID, review date, accommodation ID, plan ID, plan title, room type, room name, purpose, accompanying group, and user ratings from different aspect categories, as well as an overall score. We present statistical information about our corpus and provide insights into factors driving data drift between 2019 and 2024 using statistical approaches.
@@ 1 Introduction
Model Robustness / Practicality
@@
Intending to advance and facilitate research using customer review data in the field of Japanese NLP research and the other related fields, in this paper, we provide an up-to-date Rakuten Travel reviews corpus as shown in Table 1, ranging from the pre-COVID era to the post-COVID era. Our main contributions are summarized as follows: (1) To the best of our knowledge, in the hotel domain, no attempt has been made to release a publicly available dataset in which the collection duration is more than 15 years; (2) We distribute our corpus for easy access by any research institute for non-commercial purposes, which can be obtained via the IDR repository [^7] https://www.nii.ac.jp/dsc/idr/rakuten/; (3) We give an insight into factors of data drift in hotel review texts with statistical approaches.
@@
Table 1: Corpus Specification of Rakuten Travel Reviews Data (English translation is given in parenthesis)
@@ 3.1 Corpus Specification
This column represents a masked user name that start with prefix“user_”. Our corpus has 2,505,326 distinct users.
@@
This column refers to the posted Date of review, formatted as YYYY-MM-DD HH:MM:SS.
@@
Plan ID and Plan title
In the context of hotel reservations, a “plan” refers to a package product or campaign offered by accommodation facilities, combining lodging with associated services and benefits. It goes beyond simply “booking a room,” and is sold with various elements incorporated to meet the diverse needs and purposes of guests.
User Rating
Reviewers can rate the evaluation on a five-point scale across predefined six categories — Location, Room, Food, Bath, Services, and Amenity — as well as an overall rating. The value “5” indicates the highest evaluation, whereas a rating “1” refers to the lowest evaluation. The value “0” indicates that the user was unable to evaluate the corresponding aspect category because there was no opportunity to experience it during the stay.
@@ 3.2 Data Statistics
To verify the data draft between old and new data, we observe a gap in word distribution in reviews between 2019 and 2024, as evidenced by the chi-square test of words. We first divided a review into one or more sentences with the heuristic rules. Then, we tokenized review sentences with MeCab and the MeCab-ipadic-neologd dictionary, and counted the frequencies of unigrams, bigrams, and trigrams. We found 4,453,580 tokens with statistically significant differences at 1% level. Let us discuss factors causing data drift, especially focusing on the following factors, as indicated by the higher odds rate and chi-square value in Table 2.
@@ 4 Factors Causing Data Drift
Figure 2: Rating Distribution on aspect by aspect basis
@@
Table 2: Document frequency of words with statistically significant difference between 2019 and 2024
@@ 5 Conclusion
This paper presents an up-to-date Rakuten Travel reviews corpus to facilitate researchers at academic institutions, particularly universities, in conducting research with real-world data. To the best of our knowledge, there have been no prior attempts to publicly release hotel review data with a collection period exceeding 15 years. This dataset is available to all universities via the IDR repository for non-commercial purposes. Through statistical analysis of word distributions between 2019 and 2024, we identified key factors causing data drift in the hotel review domain. The COVID-19 pandemic fundamentally transformed the hospitality industry by introducing new vocabulary related to remote work, contactless services, and hygiene management. With the increasing importance of SDGs, customers are increasingly evaluating and providing feedback on environmentally conscious initiatives, particularly regarding amenity distribution practices. Furthermore, a surge in inbound tourism, driven by social media trends and favorable exchange rates, has significantly impacted review content. These findings underscore the importance of using up-to-date data for model training and evaluation in real-world business applications.

### v3

@@ Abstract
This paper presents a large-scale corpus of Rakuten Travel Reviews. Our collection contains 7.29 million customer reviews for 16 years, ranging from 2009 to 2024. Each record in the dataset contains the review text, its response from an accommodation, an anonymized reviewer ID, review date, accommodation ID, plan ID, plan title, room type, room name, purpose, accompanying group, and user ratings from six aspect categories, as well as an overall score. We present statistical information about our corpus and provide insights into factors driving data drift between 2019 and 2024 using statistical approaches.
@@ 1 Introduction
Model Robustness and Practicality
@@
Intending to advance and facilitate research using customer review data in the field of Japanese NLP research and the other related fields, in this paper, we provide an up-to-date Rakuten Travel reviews corpus as shown in Table 1, ranging from the pre-COVID era to the post-COVID era. Our main contributions are summarized as follows: (1) To the best of our knowledge, in the hotel domain, no attempt has been made to release a publicly available dataset in which the collection duration is more than 15 years; (2) We distribute our corpus for easy access by any research institute, particularly, university laboratories for academic research purposes, which can be obtained via the IDR repository [^7] https://www.nii.ac.jp/dsc/idr/rakuten/; (3) We give an insight into factors of data drift in hotel review texts with statistical approaches.
@@
Table 1: Corpus Specification of Rakuten Travel Reviews Data (English translation is given in parentheses)
@@ 3.1 Corpus Specification
Reviewer ID represents a masked user name that starts with prefix“user_”. Our corpus has 2,505,326 distinct users.
@@
Review date refers to the posted date of review, formatted as YYYY-MM-DD HH:MM:SS.
@@
Plan ID
In the context of hotel reservations, a “plan” refers to a campaign offered by accommodation, combining staying with associated services and benefits. It goes beyond simply “booking a room”, and is sold with various elements incorporated to meet the diverse needs and purposes of guests. There are 985,678 unique plan IDs in our corpus.
Plan Title
Plan title represents campaign contents expressed in a short sentence.
Room Type
Room type represents a pre-defined category name for a room. Our corpus has 103,256 room types.
Room Name
Room name is associated with room features such as smoking availability, room size and capacity, expressed in a short sentence.
@@
User Rating
Rating was posted by a reviewer on a five-point scale across pre-defined six aspect categories, namely, Location, Room, Food, Bath, Services, and Amenity, as well as an overall rating. In total, our corpus has seven columns for rating scores. The value “5” indicates the highest evaluation, whereas a rating “1” refers to the lowest evaluation. The value “0” indicates that the user was unable to evaluate the corresponding aspect category since there was no opportunity to experience it during the stay.
@@ 3.2 Data Statistics
To verify the data drift between old and new data, we observe a gap in word distribution in reviews between 2019 and 2024, as evidenced by the chi-square test of words. We first divided a review into one or more sentences with the heuristic rules. Then, we tokenized review sentences with MeCab and the MeCab-ipadic-neologd dictionary, and counted the frequencies of unigrams, bigrams, and trigrams. We found 4,453,580 tokens with statistically significant differences at 1% level. Let us discuss factors causing data drift, especially focusing on the following factors, as indicated by the higher odds rate and chi-square value in Table 2.
@@ 4 Factors Causing Data Drift
Figure 2: Rating Distribution on an aspect-by-aspect basis for 2019 and 2024
@@
Table 2: Document frequency of words with statistically significant difference between 2019 and 2024 (p<0.01)
@@ 5 Conclusion
This paper presents an up-to-date Rakuten Travel reviews corpus to facilitate researchers at academic institutions, particularly universities laboratories, in conducting research with real-world data. To the best of our knowledge, there have been no prior attempts to publicly release hotel review data with a collection period exceeding 15 years. Through statistical analysis of word distributions between 2019 and 2024, we identified key factors causing data drift in the hotel review domain. The COVID-19 pandemic fundamentally transformed the hospitality industry by introducing new vocabulary related to remote work, contactless services, and hygiene management. With the increasing importance of SDGs, customers are increasingly evaluating and providing feedback on environmentally conscious initiatives, particularly regarding amenity distribution practices. Furthermore, a surge in inbound tourism, driven by social media trends and favorable exchange rates, has significantly impacted review content. These findings underscore the importance of using up-to-date data for model training and evaluation in real-world business applications.
Limitations
We grant academic research institute the access of our corpus for academic research purposes.